OpenStudy (anonymous):

Did I get this Right!! Libby invested $1150 at 5.5% annual interest rate compound quarterly. At the same bank, Laura invests $1310 at 4.0% annual interest rate compounded continuously. Who has more money at the end of 10 years and how much more does she have? - Libby and she has $1985.79

OpenStudy (amistre64):

what was your process?

OpenStudy (anonymous):

I used the formula A= P(1=r/n)^nt for Libby and the formula A=Pe^rt for Laura.

OpenStudy (amistre64):

You calculated Libbys amount after 10 years .... but that is not what the question asks for

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=1150%281%2B.055%2F4%29%5E%284*10%29 libby has a balance of about: 1985.78 laura has a balance of about: 1954.26 libby has more, but how much more?

OpenStudy (anonymous):

I dont get it. They asked who had more money after ten years and how much they had. Wouldnt the answer be Libby and that she had $1985.78?

OpenStudy (amistre64):

Who has more money at the end of 10 years and how much more does she have? ^^^^^^^^^

OpenStudy (amistre64):

if you have 3 jelly beans, and someone else has 5 jellybeans how many more jelly beans does the other person have than you?

OpenStudy (anonymous):

That person has 2 more jelly beans than i do. . . .

OpenStudy (amistre64):

yeah :) libby has how much more money than luara

OpenStudy (anonymous):

I know. I'm sure i said Libby has more money than Laura.

OpenStudy (amistre64):

yeah ... she does. and how much more does she have than luara? its not 1985 more ....

OpenStudy (anonymous):

Oh is it . . . . $31.52.